OpenStudy (anonymous):

Which point is on the graph of the equation 3x -2y = 6

OpenStudy (anonymous):

@Ashleyisakitty @jigglypuff314

pooja195 (pooja195):

You need to find the X and Y intercept

pooja195 (pooja195):

Do you know how to do that?

pooja195 (pooja195):

@Nikki85

pooja195 (pooja195):

Helps find Y \[\huge~\rm~3(0) -2y = 6 \] Helps find X \[\huge~\rm~3x -2(0) = 6\]

pooja195 (pooja195):

Need help?

OpenStudy (anonymous):

I got 3,2

pooja195 (pooja195):

Try that 1st one again

OpenStudy (anonymous):

OK

pooja195 (pooja195):

And remember (x,y)

OpenStudy (anonymous):

2,3 or 4,3

pooja195 (pooja195):

Not quite \[\huge~\rm~6\div-2=?\]

OpenStudy (anonymous):

-3

pooja195 (pooja195):

Good so we have y=-3 x=2 Now you plot those on a graph

OpenStudy (anonymous):

X is 0,-3 Y is 0,2

OpenStudy (anonymous):

O

pooja195 (pooja195):

You just plot them here use this https://www.desmos.com/calculator

OpenStudy (anonymous):

Ok

pooja195 (pooja195):

Hope i helped! :)

OpenStudy (anonymous):

R u there, so once I graph it I count and I still got 2,3 after Making my slope

pooja195 (pooja195):

x-intercept: (2,0) y-intercept: (0,−3)

pooja195 (pooja195):

You dont need the slope the directions dont tell you that

OpenStudy (anonymous):

O

OpenStudy (anonymous):

Ok Thank you

pooja195 (pooja195):

No probs!:)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~To~OpenStudy!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (anonymous):

Cool thanks for the info

pooja195 (pooja195):

:)

OpenStudy (anonymous):

Let's solve for x. 3x−2y=6 Step 1: Add 2y to both sides. 3x−2y+2y=6+2y 3x=2y+6 Step 2: Divide both sides by 3. 3x/3=2y+6/3 x=2/3y+2

OpenStudy (anonymous):

Answer is x=2/3y+2

pooja195 (pooja195):

???